How do I calculate;
I have 2 wells, well A produces water with 18 ppb of a mineral
Well B produces water with 3 ppb of the same mineral.
How many gallons per minute should I blend of well A water with how many gallons per minute of well B water to produce a combined water with a 5 ppb concentration when I need to pump 2400 Gallons per minute total flow from both wells?

Answer by josmiceli(19441)

Let A equal the number of gallons pumped from well A in one minute to satisfy the conditions of the problem.
Let B equal the number of gallons pumped from well B in one minute to satisfy the conditions of the problem.
I need to pump 2400 Gallons per minute total flow from both wells.
A + B = gallons pumped from both wells in one minute, combining the first two conditions
A + B = 2400, combining all three conditions
change ppb to k / gal or kpg
A(18 kpg) + B(3 kpg) = 2400(5 kpg)
A(18 kpg) + (2400 - A)(3 kpg) = 12000 kpg
the units, kpg, all cancel
15A + 7200 -3A = 12000
12A = 4800
A = 400 gal/min
B = 2000 gal/min